# Overcoming the Limitations of Layer Synchronization in Spiking Neural Networks

### v1

@@ Abstract
Currently, neural-network processing in machine learning applications relies on layer synchronization, whereby neurons in a layer aggregate incoming currents from all neurons in the preceding layer, before evaluating their activation function. This is practiced even in artificial Spiking Neural Networks (SNNs), which are touted as consistent with neurobiology, in spite of processing in the brain being, in fact asynchronous. A truly asynchronous system however would allow all neurons to evaluate concurrently their threshold and emit spikes upon receiving any presynaptic current.
Omitting layer synchronization is potentially beneficial, for latency and energy efficiency, but asynchronous execution of models previously trained with layer synchronization may entail a mismatch in network dynamics and performance.
We present a study that documents and quantifies this problem in three datasets on our simulation environment that implements network asynchrony, and we show that models trained with layer synchronization either perform sub-optimally in absence of the synchronization, or they will fail to benefit from any energy and latency reduction, when such a mechanism is in place.
We then “make ends meet” and address the problem with unlayered backprop, a novel backpropagation-based training method, for learning models suitable for asynchronous processing. We train with it models that use different neuron execution scheduling strategies, and we show that although their neurons are more reactive, these models consistently exhibit lower overall spike density (up to 50%), reach a correct decision faster (up to 2x) without integrating all spikes, and achieve superior accuracy (up to $\sim$10higher). Our findings suggest that asynchronous event-based (neuromorphic) AI computing is indeed more efficient, but we need to seriously rethink how we train our SNN models, to benefit from it.
@@ 1 Introduction
Artificial Neural Networks (ANNs) are the foundation behind many of the recently successful developments in AI, such as in computer vision [54, 58] and natural language processing [57, 5]. To match the complexity of the ever more demanding tasks, networks have grown in size, with advanced large language models having billions of parameters [65]. With this the power consumption exploded [28], limiting the deployment to large data centers. In an effort to learn from our brain’s superior power efficiency, and motivated in neuroscience research, SNNs [30] bolster as an alternative. They use discrete binary or graded spikes (events) for communication, are suited for processing sparse features [18], and when combined with asynchronous event-based processing are assumed to enhance latency and energy efficiency (and scalability). Sparsity leads to fewer synaptic operations, resulting in low energy consumption, and asynchronous operation potentiates concurrent evaluation of all neurons in the network purely event-driven, leading to low latency.
Conventional highly parallel ANN compute accelerators, such as Graphics Processing Units (GPUs) and Tensor Processing Units (TPUs), which are primarily optimized for dense vectorized matrix operations, face inherent challenges in exploiting sparsity for improving their energy efficiency. Targeting the commonplace practiced way of executing ANNs layer-after-layer has left them with poor support for asynchronous processing too (for improving latency). At best, they parallelize processing within a layer and/or pipeline processing across layers. This leaves an exploration space for neuromorphic processors that try to excel in handling the event-driven nature of SNNs and leverage asynchronous concurrent processing, offering efficiency advantages in various tasks [19, 21, 36].
However, despite this advancement, the training of SNNs today very often conveniently relies on conventional end-to-end ANN training methods for performance [9], which organize/synchronize computations per-layer rather than event-driven per neuron [17]. Specifically, at any given discrete timestep within a neuron layer, first, the total of all presynaptic currents (from the preceding layer) must be computed and integrated, before postsynaptic neurons in the current layer update their state and evaluate their activation function (i.e. emitting new spikes). For consistency, neuron evaluation in one layer must thus complete and synchronize before proceeding to evaluate neurons of a next layer. This breadth-first processing approach (with per layer synchronization), while it facilitates use of vectorized computing hardware (such as GPUs) during training, it introduces dependencies on synchronization that could impact model accuracy if altered during inference. Consequently, even asynchronous neuromorphic processors, such as Loihi [10], have integrated mechanisms to ensure (and enforce) layer synchronization.
This leaves a crucial (efficiency) aspect of SNNs relatively unexplored: the ability to allow spiking completely asynchronously across the network without having separate phases for integration and firing, just like in our brain [63]. In that neurophysical modus operandi, neurons can fire and receive currents anywhere in the network at any time, completely independent, a concept we term "network asynchrony". Allowing network asynchrony can be advantageous [42], as we confirm in our results. Spike activity can quickly propagate deep into the network without being bound by synchronization barriers, reducing latency. Furthermore, adhering to layer synchronization could lead to increased computational overhead as the network scales, as suggested by Amdahl’s law [46]. This implies that the overhead grows non-linearly by adding more computational units to a group that needs to be synchronized at some point in time [60]. With network asynchrony, such groups can be kept smaller, and the number of synchronization moments can be minimized, potentially reducing the waiting time.
@@
In this paper, we demonstrate this problem and explore solutions. Using a simulator that implements the concept of network asynchrony, we provide quantitative results based on 3 benchmark datasets (with different spatio-temporal structures) that show the performance degradation and latency/energy inefficiency resulting from changes in model dynamics when trained with layer synchronization and later deployed for asynchronous inference. We then explore potential solutions by proposing a novel gradient (backpropagation-based) training method, that can be parameterized with various neuron execution scheduling strategies for asynchronous processing, and vectorization abilities that may be available in different neuromorphic processors. We show that using this training method, it is possible to prepare models that achieve superior accuracy and, at the same time, can save energy and inference latency under asynchronous processing (when compared to the conventional breadth-first processing in GPUs). This work opens a space with unique potential for design space explorations aimed to bridge the efficiency gap between neural network model training and asynchronous processor design.
@@ 2 Related work
The term “asynchronous processing” is often used whenever neurons can be active in parallel and communicate asynchronously, even if some synchronization protocol is enforced to control the order of spikes and synaptic current processing [44, 22, 51, 61]. Another kind of asynchrony is related to input coding[15]. In this context, synchronizing means grouping spikes (events) into frames, a topic that has extensively been researched [18, 33, 43, 45, 56]. Neither of the two, however, is the focus of this work. Here, asynchrony refers solely to activity within the network not being artificially bound by any order restriction. This has been researched for simulating biologically accurate neural networks [29, 32, 31], but remains under-explored in the context of executing SNNs for machine learning. In this context, grouping and processing in layers is just one convenient way [34].
Few event-driven neuromorphic processors, such as $\upmu$Brain [53] and Speck [6], are fully event-driven and lack any layer synchronization mechanism. This makes them notorious for training out-of-the-box with mainstream model training tools that rely on per-layer synchronization (e.g., PyTorch). Speck developers propose to train their models with hardware in-the-loop [27] to reduce the mismatch between the training algorithms and the inference hardware. However this method does not provide a general solution for training asynchronous neural networks. Others, like SpiNNaker [14] and TrueNorth [1], use timer-based synchronization, or Loihi [10] and Seneca [55] use barrier synchronization between layers, in order to warrant that the asynchronous processing dynamics on hardware are aligned at layer boundaries with models trained in software (with layer synchronization). This, however, entails an efficiency penalty, as we will show.
@@
Functionally, the most suitable type of model for end-to-end asynchronous processing is probably rate-coded SNN models, whereby neurons can integrate state and communicate independently from any other neuron. These models are trained like ANNs [64, 26, 11] or converted from pre-trained ANNs [47, 25]. Therefore, to be accurate under asynchronous processing, it is required to run the inference for a long time, reducing the latency and energy benefit of using SNNs [49].
@@
Alternative to rate-coding models are temporal-coding models, with time-to-first spike (TTFS) [41, 52, 23, 7] or order encoding [4]. They are very sparse (hence energy efficient) but very cumbersome and elaborate to convert from ANNs [39, 52] or train directly [23, 7], less tolerant to noise [7], and their execution so far, while event-based, requires some form of synchronization or a reference time. Efficiency is thus only attributable to the reduced number of spikes, all of which need to be evaluated in order before a decision is reached. Other alternative encodings for event-based processing of SNNs include phase-coding [24] and burst coding [40], which are, however, no more economical than rate coding and have not been shown, to our knowledge, to attain competitive performance.
@@
The approach presented in this paper is, in fact, unique in enabling the trainability of models for event-based asynchronous execution, providing efficiency from processing only a subset of spikes, and delivering consistent performance and tolerance to noise. Also relevant to the works in this paper are SparseProp [12] and EventProp [59] on efficient event-based simulation and training, respectively. EventProp[59] is potentially more economical than discrete-time backpropagation for training event-based models for asynchronous processing (and fully compatible with the work in this paper), but it has not been shown how its complexity scales beyond 1 hidden layer. SparseProp [12] proposes an efficient neuron execution scheduling strategy for asynchronous processing in software simulation. An effective hardware implementation of this scheduling strategy can be found in [35], which has inspired the herein proposed “momentum schedule”. Moreover, we advance this research by demonstrating how to train SNN models using this event scheduler.
@@ 3 Methods for simulating and training of asynchronous SNNs
In this section, we provide an overview of our methodology. We first introduce our simulator for SNN processing with network asynchrony and then show how this is used for training.
@@ 3.1 Simulating asynchronous SNNs
The SNNs used in this work consist of $L$ layers of Leaky Integrate and Fire (LIF) neurons [18], with each layer $l$ for $1\leq l\leq L$ having $N^{(l)}$ neurons and being fully-connected to the next layer $l+1$, except for the output layer. Each input feature is connected to all neurons in the first layer. Every connection has a synaptic weight. Using these weights, we can compute the incoming current $x$ per neuron resulting from spikes, as per equation 3.
@@
Each LIF neuron has a membrane potential exponentially decaying over time based on some membrane time constant $\tau_{m}$. We use the analytical solution (see equation 4) to compute the decay. By keeping track of the membrane potential $u[t]$ and the elapsed time $\Delta t$ since time $t$, it is possible to precisely calculate the membrane potential for $t+\Delta t$. Therefore, computations are required only when $x[t+\Delta t]>0$.
@@
To determine if a neuron spikes, a threshold function $\Theta$ checks if the membrane potential exceeds a threshold $U_{\text{thr}}$, following equation 5. If $\Theta(u)=1$, the membrane potential is immediately hard reset to 00. Soft reset is another option in the simulator often defended in the literature [16] as well as refractoriness [48]. However for simplicity here we only experiment with hard reset.
3.1.1 Event-driven state updates
@@
We can vectorize the computations introduced in the previous section to perform them on a per-layer basis. This is the "layered" inference approach, which implies layer synchronization, since all neurons of one layer evaluate their state at the same instant. An alternative to this is an event-driven approach, where the computations for state updates are applied in response to new spike arrivals. Spikes can be emitted at any time by any neuron, affecting the states of postsynaptic neurons independently of others. Such an approach can achieve a true representation of network asynchrony.
When using event-driven updates, the network dynamics evolve with each spike. A single spike can generate multiple currents, each linked to a specific time $t$, determined by the initiating input activity. Let $K_{t}$ represent the atomic computational unit that is executed for an input current received at time $t$ and that will update the state of a single neuron and evaluate its threshold function (activation). The exact computations for a LIF neuron can be found in section A.2. The order in which $K_{t}$ is executed across neurons inside the entire network (and not just in one layer) affects the output of the network due to the non-linear nature of $K_{t}$ (see figure 1 for an example of the effect this can have). For simplicity, it is assumed that the currents resulting from the same spike are processed in a group at once. This processing is done when an emitted spike is selected for processing, i.e., when the spike is propagated. This may not necessarily be the same moment as when the spike was emitted due to asynchrony. Because of these assumptions, analyzing the spike propagation order gives the same insights as analyzing the execution order of $K_{t}$.
@@
Unlike the layered approach, with event-driven state updates neurons can theoretically fire more than once at the same time $t$. We prevent this by inducing a neuron to enter a refractory state for the remainder of time $t$. In this state, the neuron keeps its membrane potential at 0. This makes the model causally simpler, more tractable, and more "economic" when energy consumption is coupled to spike communication. These aspects also give an explicit retrospectively justification for the neurophysically observed refractoriness [3].
Figure 1: With network asynchrony, currents from neurons $x_{i}$ in layer $l$ may come to neuron $y$ in layer $l+1$ at any time in any order. It may spike before all inputs have been received in a discrete timestep. Illustrated is the exceedance of the firing threshold at point $A$. All inputs, and thus information after $A$ is missed, i.e., the spiking decision is made based on partial information. In this example, if layer synchronization is enforced, neuron $y$ would wait for 100% of the input and will not spike.
The concept of time
We apply input framing in discrete timesteps as typically done for inputs from a DVS sensor. The timestep size refers to the duration within which all input spikes are collected into a single matrix (also called frame), before being propagated throughout the network collectively in what is termed a forward pass. This input framing process precedes any network activity and follows the completion of all activity from the last forward pass. The timestamps of input spikes are rounded down to the nearest multiple of the timestep size.
The timestamps on the input spikes provide a meaningful measure of time. Upon arrival, they initiate network activity, making all new spikes contingent on their timing. Thus, they serve as the markers for the progression of time in timesteps, while the order of spike propagation though the network establishes another concept of “time” within a forward pass (or “micro-time” between timesteps).

### v2

# Exploring the Limitations of Layer Synchronization in Spiking Neural Networks

@@ Abstract
Neural-network processing in machine learning applications relies on layer synchronization. This is practiced even in artificial Spiking Neural Networks (SNNs), which are touted as consistent with neurobiology, in spite of processing in the brain being in fact asynchronous. A truly asynchronous system however would allow all neurons to evaluate concurrently their threshold and emit spikes upon receiving any presynaptic current. Omitting layer synchronization is potentially beneficial, for latency and energy efficiency, but asynchronous execution of models previously trained with layer synchronization may entail a mismatch in network dynamics and performance. We present and quantify this problem, and show that models trained with layer synchronization either perform poorly in absence of the synchronization, or fail to benefit from any energy and latency reduction, when such a mechanism is in place. We then explore a potential solution direction, based on a generalization of backpropagation-based training that integrates knowledge about an asynchronous execution scheduling strategy, for learning models suitable for asynchronous processing. We experiment with 2 asynchronous neuron execution scheduling strategies in datasets that encode spatial and temporal information, and we show the potential of asynchronous processing to use less spikes (up to 50%), complete inference faster (up to 2x), and achieve competitive or even better accuracy (up to $\sim$10% higher). Our exploration affirms that asynchronous event-based AI processing can be indeed more efficient, but we need to rethink how we train our SNN models to benefit from it. (Source code available at: https://github.com/RoelMK/asynctorch)
@@ 1 Introduction
Artificial Neural Networks (ANNs) are the foundation behind many of the recently successful developments in AI, such as in computer vision Szegedy et al. (2017); Voulodimos et al. (2018) and natural language processing Vaswani et al. (2017); Brown et al. (2020). To match the complexity of the ever more demanding tasks, networks have grown in size, with advanced large language models having billions of parameters Zhao et al. (2024). With this, the power consumption has exploded Luccioni et al. (2023), limiting the deployment to large data centers. In an effort to learn from our brain’s superior power efficiency, and motivated by neuroscience research, SNNs Maass (1997) bolster as an alternative. They use discrete binary or graded spikes (events) for communication, are suited for processing sparse features He et al. (2020), and when combined with event-based asynchronous processing are assumed to reduce latency and improve energy efficiencyZeki (2015). Sparsity can save synaptic operations, thus lowering energy consumption, and asynchronous operation enables concurrent evaluation of all neurons anywhere in the network purely event-driven, leading to low latency.
Figure 1: (A) Event-based processing. (B) Event-based processing with per-layer synchronization: latency is a function of the events in the system as well as the number of layers (synchronization barriers). (C) Asynchronous (end-to-end) event-based processing: event processing does not encounter synchronization barriers, and so inference latency is determined solely by the subset of the total number of events processed until a decision is made at the output layer. Color codes indicate at which layer events were fired. The x-axis of the plots represents time/order that events were processed. The vertical planes in (B) represent layer-synchronization barriers (also indicating bottlenecks in network routing and memory I/O). The difference between (B) and (C) can be understood through an analogy: bikes racing in a circuit versus on a public road with a sequence of traffic lights.
In an event-based processing system (figure 1.A), a spike or event in an SNN (or non-zero activation in an ANN) can independently and locally trigger some downstream computations (that may or may not have a global effect). Event-based processing manifests essentially by comparison to vector-based processing, which groups (activation) computations in batches.
By asynchronous processing we understand the fact that spikes (or activations) can be processed at any time (or in any order), and also anywhere in the network, independently of any other spikes and unhindered by artificial conditions that delay their immediate propagation (such as waiting for all other neurons in the same layer to integrate their currents or evaluate their thresholds). Typically, for a large network of neurons only the approximate timing of events plays a role in asynchrony London et al. (2010), and the resulting order is in often the main information-carrying signal (according to neuroscience literature that underpins order codings) Thorpe et al. (2001). Approximate or relative timing makes asynchronous processing viable in digital accelerators even though they discretize time. Asynchronous processing, as perceived in this work, is illustrated (and visually represented through an analogy) in figure 1.C
It needs to be meticulously clarified that asynchronous processing within the network, and at the hardware processing system level (which is the focus of this work), is distinct from what is customarily referred to as asynchrony at the model input. Specifically, at the model level in RNNs, we colloquially refer to timesteps as the discretization of continuous external stimuli into a sequence of consecutive steps for input to the network model. Asynchrony in the latter sense (as commonly seen in SNN literature) refers to the fact that the stimulus is sampled across the input dimension and can be presented to the model not in a single step, but rather distributed i.i.d. across several such discrete timesteps. To limit ambiguity, we call this input asynchrony  [^1] Notably a few approaches, such as in Zhu et al. (2022) quote as asynchrony the manipulation through learning of the shape and distribution of stimulus (input spikes) across timesteps with the goal of improving model performance (accuracy), assuming however execution with layer-synchronized processing (event-based or vectorized alike)., and we distinguish it from (in-)network asynchrony which we study in this paper, namely the relative propagation of spatial signals inside the network within individual timesteps.  [^2] As will be seen later, we also need a notion of discretization of (processing) steps through the network, which can be variable and decoupled from the number of layers. We refer to these steps as forward steps. Thus, while timesteps discretize processing (in complete forward model evaluations) across stimulus presentations, forward steps discretize processing within a timestep from the onset of (one timestep’s) stimulus presentation until the output neurons conclude the model evalution for that timestep. Moreover, with input asynchrony, by common convention, processing of the stimulus in one timestep inside the network has been assumed until now to be as taking place "atomically" through a deterministic process and sequence: breadth-first across neurons one-layer-after-the-other. This convention serves the purpose that spikes (activation state) from two distinct timesteps, but also between layers, do not interfere with one another. Forward inference in two timesteps interacts only through "neuron state caching" in each layer, by means of (not resetting) the membrane voltages (hidden state in RNNs). But, from the viewpoint of this work, this type of processing inside the network is still synchronous (despite input asynchrony) as depicted in figure 1.B. Moreover, we argue this per-layer synchronization dulls the dynamics of the stimulus (data-generating process); is not quite what happens in the brain (SNNs); and neither is the type of processing neuromorphic processors are primarily aimed for.
Asynchronous processing in the brain Zeki (2015) has been connected to energy and latency efficiency because information dynamics can express freely and propagate signals promptly through neuron clusters triggering early outputs and leveraging fast decisions even before other neuron clusters of slower dynamics have finished processing the input stimulus (and generated downstream activations). Specifically, there is no known process in the brain that consistently and uniformly forces neurons in one layer (or column) to finish their computations before the neurons in another downstream layer assume processing of their input currents. On the contrary, the more intensely the stimulus manifests in some regions (subgroups of neurons), the faster dynamics are reinforced, leading to a non-uniform (unaligned) and non-synchronized propagation of spikes across regions. These dynamics are also likely modulating and/or are modulated along certain pathways by heterogeneity in axonal or synaptic delays and different state decay processes across neurons.
In other words, through asynchronous processing, in-network expressed dynamics should lead to computation economy. The work in this paper puts this hypothesis to test. It presents a first of its kind study of how current synchronous training impacts such an asynchronous modus operandi, how can we possibly train good performing robust models for asynchronous processing, and if energy and latency efficiency stem from such a modus operandi.
@@
Using a simulation environment that implements the concept of network asynchrony and can emulate the processing behavior of a number of neuromorphic accelerators, we provide quantitative results on benchmark datasets (with different spatio-temporal information content) and network topologies of two or more hidden layers, which show performance degradation and latency/energy inefficiency resulting from changes in model dynamics when trained with layer synchronization and later deployed for asynchronous inference (processing). Next, we explore a potentially promising solution by proposing a generalization of gradient (backpropagation-based) training, that can be parameterized with various neuron execution scheduling strategies for asynchronous processing, and vectorization abilities present in various neuromorphic processors. We show that using this training method, it is possible not only to recover the compromised accuracy, but also to fulfil the expectation of saving energy and improving latency under asynchronous processing (when compared to the conventional breadth-first processing in GPUs). This work opens a path for design space explorations aimed to bridge the efficiency gap between neural network model training and asynchronous processor design.
@@ 2 Related work
Conventional highly parallel ANN compute accelerators, such as Graphics Processing Units (GPUs) and Tensor Processing Units (TPUs), which are primarily optimized for dense vectorized matrix operations, face inherent challenges in exploiting unstructured and temporal sparsity for improving their energy efficiency. Targeting the common practice of executing ANNs with per-layer synchronization bottlenecks has left them with poor support for asynchronous processing (for improving latency or energy consumption). At best, they parallelize processing within a layer and/or pipeline processing across layers. This leaves an exploration space for neuromorphic processors that try to excel in handling the event-driven nature of SNNs and leverage asynchronous concurrent processing, offering efficiency advantages in various tasks.
However, the training of SNNs today more and more often relies conveniently, for performance, on conventional end-to-end ANN training methods Dampfhoffer et al. (2023), based on synchronizing computations per-layer rather than event-driven per neuron Guo et al. (2023a). First, the total of all presynaptic currents from the preceding layer must be computed and integrated before postsynaptic neurons in the current layer update their state and evaluate their activation function (i.e. emit new spikes). For consistency, neuron evaluation in one layer must thus complete and synchronize before proceeding to evaluate neurons of the next layer.
[…]
@@
Functionally, the most suitable type of model for end-to-end asynchronous processing is probably rate-coded SNN models, whereby neurons can integrate state and communicate independently from any other neuron. These models are trained like ANNs Zenke & Vogels (2021); Lin et al. (2018); Diehl et al. (2015) or converted from pre-trained ANNs Rueckauer et al. (2017); Kim et al. (2020). Therefore, to be accurate under asynchronous processing, it is required to run the inference for a long time, reducing the latency and energy benefit of using SNNs Sengupta et al. (2019).
@@
Alternative to rate-coding models are temporal-coding models, with time-to-first spike (TTFS) Park et al. (2020); Srivatsa et al. (2020); Kheradpisheh & Masquelier (2020); Comşa et al. (2022) or order encoding Bonilla et al. (2022). They are very sparse (hence energy efficient) but very cumbersome and elaborate to convert from ANNs Painkras et al. (2013); Srivatsa et al. (2020) or train directly Kheradpisheh & Masquelier (2020); Comşa et al. (2022), less tolerant to noise Comşa et al. (2022), and their execution so far, while event-based, requires some form of synchronization or a reference time. Efficiency is thus only attributable to the reduced number of spikes, all of which need to be evaluated in order before a decision is reached. Other alternative encodings for event-based processing of SNNs include phase-coding Kim et al. (2018) and burst coding Park et al. (2019), which are, however, no more economical than rate coding and have not been shown, to our knowledge, to attain competitive performance.
@@
The approach presented in this paper is, in fact, unique in enabling the trainability of models for event-based asynchronous execution, providing efficiency from processing only a subset of spikes, and delivering consistent performance and tolerance to noise. Also relevant to the works in this paper are SparseProp Engelken (2024) and EventProp Wunderlich & Pehle (2021b) on efficient event-based simulation and training, respectively. EventProp Wunderlich & Pehle (2021b) is potentially more economical than discrete-time backpropagation for training event-based models for asynchronous processing (and fully compatible with the work in this paper), but it has not been shown how its complexity scales beyond 1-2 hidden layers. SparseProp Engelken (2024) proposes an efficient neuron execution scheduling strategy for asynchronous processing in software simulation. An effective hardware implementation of this scheduling strategy can be found in Monti et al. (2017), which has inspired the herein proposed “momentum schedule”. Moreover, we advance this research by demonstrating how to train SNN models using this event scheduler.
the fact that neurons evaluate their threshold immediately upon receiving a synaptic input (spike) from each fan-in synapse independently (as opposed to aggregating all synaptic inputs and the evaluating the threshold once – where they can often cancel each other out); the relaxation of forcing synchronization of processing at every layer; and the adoption of dynamic scheduling of neuron execution across the entire network which continuously adapts (stochastically) to flow dynamics of stimulus/activations or neuron membrane states. The last last two “ingredients” allow neurons deep inside the network to fire before processing at early layers has finished (if the flow dynamics encourage it).
@@ 3 Methods for simulating and training of asynchronous SNNs
In this section, we detail our methodology. First, we detail (section 3.1) how we simulate the unconstrained behavior of various (digital) neuromorphic event-based accelerators capable of asynchronous SNN processing inside the neural network (i.e., implement network asynchrony), by leveraging the following principles: (a) allow neurons to evaluate their thresholds independently for every received synaptic current; (b) abolish the constraint to synchronize activations at each layer; (c) adopt dynamic scheduling of neuron execution across the entire network that continuously adapts to the flow dynamics of activations/stimulus or relative neuron states; and (d) support limited-length vectorization. (Details about the accelerators we considered, and the parameterizability of the simulator to support them, are given in the tables in appendix A.5)
Then in section 3.2 we explain how these principles can be applied to backpropagation training (leading to an approach we call unlayered backpropagation) to prepare good performing, robust and efficient asynchronous models.
@@ 3.1 Simulating asynchronous SNNs
The SNNs used in this work consist of $L$ layers of Leaky Integrate-and-Fire (LIF) neurons He et al. (2020), with each layer $l$ for $1\leq l\leq L$ having $N^{(l)}$ neurons and being fully-connected to the next layer $l+1$, except for the output layer. Each input feature is connected to all neurons in the first layer. Every connection has a synaptic weight. Using these weights, we can compute the incoming current $x$ per neuron resulting from spikes, as per equation 3 in appendix A.1.
@@
Each LIF neuron has a membrane potential exponentially decaying over time based on some membrane time constant $\tau_{m}$. We use the analytical solution (see equation 4 in appendix A.1) to compute the decay. By keeping track of the membrane potential $u[t]$ and the elapsed time $\Delta t$ since time $t$, it is possible to precisely calculate the membrane potential for $t+\Delta t$. Therefore, computations are required only when $x[t+\Delta t]>0$.
@@
To determine if a neuron spikes, a threshold function $\Theta$ checks if the membrane potential exceeds a threshold $U_{\text{thr}}$, following equation 5 in appendix A.1. If $\Theta(u)=1$, the membrane potential is hard reset to 0. Soft reset is another option Guo et al. (2022) as well as refractoriness Sanaullah et al. (2023). However, for simplicity here we only experiment with hard reset.
3.1.1 Asynchronous inference with event-driven state updates
@@
We can vectorize the computations introduced in the previous section on a per-layer basis. This is herein referred to as the "layered" inference approach, which implies layer synchronization, since all neurons of one layer need to evaluate their state, before any neuron in a down-stream layer does the same. An alternative to this situation is an event-driven approach, where the computations for state updates are applied in response to new spike arrivals. If additionally spikes can be emitted (threshold evaluation) at any point in time (e.g in response to any individual synaptic current), by any neuron in the network, affecting the states of postsynaptic neurons independently of others, then (we assume) this approach can achieve a true representation of network asynchrony.
When using network asynchrony, the network dynamics evolve with each spike. Any single spike can generate multiple currents downstream, linked to a specific stimuli at timestep $t$, determined by the initiating input activity. Let $K_{t}$ represent the atomic computation that is executed at one neuron in the network, in response to input synaptic currents and triggered by stimulus presented at the network at timestep t, and which will update the state of the neuron and evaluate its threshold function (activation). The exact computations for a LIF neuron can be found in section A.2.
Notice that whether $K_{t}$ executes in response to input current from each synapse independently, or in response to the aggregate synaptic current from all fan-in synapses collectively, leads to different activation behavior/patterns and hence neuron dynamics (see figure 2). Also, the relative order in which $K_{t}$ executes among neurons across the entire network (as opposed to within just one layer – due to synchronization) affects the output of the network due to the non-linearity entailed in $K_{t}$. Here, for simplicity, it is assumed that the fan-out currents resulting from a spike generation are propagated as a group at once (e.g. no heterogeneous propagation delays exist). Because of network asynchrony however the recipient neurons will not necessarily process them simultaneously or in fixed order, but rather opportunistically as dictated by a (dynamic) scheduling algorithm. This dynamic scheduling captures the asynchronous hardware behaviour/capabilities (more about this is given in section 3.1.2). Because of these assumptions, analyzing the spike propagation order gives the same insights as analyzing the execution order of $K_{t}$ (across neurons in the network).
@@
Overall, unlike the layered approach, in this modus-operandi because of the per-input current evaluation of the firing threshold neurons can fire more than once at the same timestep $t$ (in response to the same input stimulus). We can prevent this by inducing a neuron to enter a refractory state for the remainder of time $t$. In this state, the neuron keeps its membrane potential at 0. This makes the model causally simpler, more tractable, and more "economic" when energy consumption is coupled to spike communication. These aspects may also give explicit retrospective justification for the neurophysically observed refractoriness Berry & Meister (1997). However, the observations we report in the results hold true both when we allowed neurons in the network to enter refractoriness after firing a spike, as well as when there is no refractoriness.
Figure 2: Layer synchronization ensures that activations from all neurons ($x_{i}$) in a layer are available simultaneously to any post-synaptic neuron ($y$) in the next layer. Hence they are integrated together and the threshold of the post-synaptic neuron is evaluated only once (see top-left graph). As a result the currents arriving from different synapses often cancel each other out given opposite sign weights, resulting in fewer spikes and a different firing pattern than when each synaptic current is processed independently (see bottom-left graph), where each input is more likely to trigger a spike, as is the case in asynchronous processing. Thus, in SNNs, layer synchronization leads to the volume of spikes being reduced and the activation flow dynamics being dulled. This is not an issue (at least not as grave) in ANNs because activations communicate relative magnitudes.
Two resolutions of time
We apply input framing as typically done for inputs from a DVS sensor, accumulating timestamped events in discrete time-bins, and providing these frames as inputs to the network in discrete timesteps. The timestep size refers to the time-interval between the timestamps of the first (or last) events in the time-bins of two consecutive timesteps. Timesteps correspond to one concept of timing (resolution) with same semantics as in RNNs, and which we refer to here as macro-time. At each timestep, input event frames initiate network activity, in a so-called forward pass, making all new spikes contingent on the timestep.
The order of spikes propagated though the network during a timestep establishes another notion of “time” in a forward pass (we call it “micro-time”), which is discretized in so-called forward steps. A forward step is associated with the processing of one or a group of events (in case of vectorization), and the number of forward steps measures time to complete inference. One can realize that in case of layered inference the number of forward steps is fixed and equals the number of layers, but this is not the case for asynchronous processing.